(3/x)+(x/3)=(5/2)
How would I solve this??

Found 2 solutions by TimothyLamb, MathTherapy:
Answer by TimothyLamb(4379)

(3/x) + (x/3) = (5/2)
(3/x)x + (x/3)x = (5/2)x
3 + (1/3)x^2 = (5/2)x
(1/3)x^2 + (-5/2)x + 3 = 0
---
the above quadratic equation is in standard form, with a=1/3=0.3333333, b=2.5, and c=3
---
to solve the quadratic equation, plug this:
0.3333333 2.5 3
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the vertex of the quadratic equation is a minimum point at: ( -3.75000038, -1.68750047 )
---
Answer 2:
the quadratic equation has two real roots:
root1 = -1.49999995
root2 = -6.0000008
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by MathTherapy(10552)


(3/x)+(x/3)=(5/2)
How would I solve this??

DUPLICATE!!